OpenStudy (anonymous):

write a logarithmic form for 2^7=128

OpenStudy (anonymous):

if \[a^b \]=N then\[\log_{a}N\]=b

OpenStudy (anonymous):

plug that in

OpenStudy (anonymous):

log2^128=7

OpenStudy (anonymous):

I guess i dont really understand

OpenStudy (anonymous):

here's something that may help you understand, what you are trying to find is what number do you have to raise 2 to, to get 128 in this case is the number 7

OpenStudy (anonymous):

I get that part but how do I right it in logarithmic form ?

OpenStudy (anonymous):

I thought i just told you

OpenStudy (anonymous):

think of "log" as another word for power, so Log of base 2 raised to 128 = 7, or log2^128=7

OpenStudy (anonymous):

ok that helps sorry takes me a little to understand math isnt my best subject

OpenStudy (anonymous):

its alright not my best either i get some stuff easier than others

OpenStudy (anonymous):

just keep working at it, and Youtube is usually helpful for me too

OpenStudy (anonymous):

mtipton is right lol youtube taught me calc